OpenStudy (anonymous):

What is the slope of the line 2x - 5y = -10? -5/2 5/2 2/5 -2/5 none of the above

OpenStudy (anonymous):

First you have to set it up in y=mx+b form

OpenStudy (saifoo.khan):

Solve for y. You will get something like: y = mx + b After comparing you will see that "m" is the slope.

OpenStudy (anonymous):

Slope is given by negative of x coefficient divided by y coefficient...

OpenStudy (anonymous):

is equal to c

OpenStudy (anonymous):

Where??

OpenStudy (anonymous):

2/5

OpenStudy (anonymous):

im righ or wrong

OpenStudy (anonymous):

plz

OpenStudy (anonymous):

im pretty sure your right because -2/-5 id 2/5 so yes your right

OpenStudy (anonymous):

Absolutely right..

OpenStudy (anonymous):

Remember when you are given with one equation of Line: That is : \(ax + by = c\) Then, \[Slope = \frac{- (coefficient \; of \;x)}{Coefficient \; of \; y}\]

OpenStudy (anonymous):

thank u

OpenStudy (anonymous):

no problem